# A generalized e-value feature detection method with FDR control at multiple resolutions

### v4

@@ Abstract
Multiple resolutions arise across a range of explanatory features due to domain-specific structures, leading to the formation of feature groups. It follows that the simultaneous detection of significant features and groups aimed at a specific response with false discovery rate (FDR) control stands as a crucial issue, such as the spatial genome-wide association studies. Nevertheless, existing methods such as the multilayer knockoff filter (MKF) generally require a uniform detection approach across resolutions to achieve multilayer FDR control, which can be not powerful or even not applicable in several settings. To fix this issue effectively, this article develops a novel method of stabilized flexible e-filter procedure (SFEFP), by constructing unified generalized e-values, developing a generalized e-filter, and adopting a stabilization treatment. This method flexibly incorporates a wide variety of base detection procedures that operate effectively across different resolutions to provide stable and consistent results, while controlling the false discovery rate at multiple resolutions simultaneously. Furthermore, we investigate the statistical theories of the SFEFP, encompassing multilayer FDR control and stability guarantee. We develop several examples for SFEFP such as eDS-filter and eDS+gKF-filter. Simulation studies demonstrate that the eDS-filter effectively controls FDR at multiple resolutions while either maintaining or enhancing power compared to MKF. The superiority of the eDS-filter is also demonstrated through the analysis of HIV mutation data.
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 rigorously frames the problem of feature detection with multilayer FDR control and introduces necessary preliminaries. Section 3 introduces the proposed SFEFP and relevant theoretical results. In Section 4, we propose eDS-filter and eDS+gKF-filter as examples of SFEFP. Other applications and possible extensions of SFEFP are also discussed. Section 5 presents the performance of our proposed methods on simulated data. In Section 6, the eDS+gKF-filter is applied to analyze HIV mutation data. Section 7 discusses the applicable scenarios and the future work of SFEFP. The proofs of all propositions, theorems, and lemmas in the subsequent sections of the paper are provided in the supplementary material.
@@ 3.1 FEFP: Flexible E-Filter Procedure
until until $t^{(1)},\ldots,t^{(M)}$ are all unchanged;
@@ 3.2 SFEFP: Stabilized Flexible E-Filter Procedure
for $r=1,\ldots,R$ do
   Separately process any fixed detection procedure $\mathcal{G}^{(m)}_{r}\in\mathcal{K}_{\text{finite}}\bigcup\mathcal{K}_{\text{asy}}$ with level $\alpha_{0}^{(m)}$, $m\in[M]$.
  Compute the generalized e-value $e_{gr}^{(m)}=\frac{G^{(m)}}{\widehat{V}_{\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})}\bigvee\alpha_{0}^{(m)}}\cdot\mathbb{I}\left\{g\in\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})\right\},~g\in[G^{(m)}],~m\in[M]$.
@@
Compute the averaged generalized e-value $\overline{e}_{g}^{(m)}=\frac{1}{R}\sum\limits_{r\in[R]}w^{(m)}_{r}e_{gr}^{(m)},~g\in[G^{(m)}],~m\in[M].$
@@
     where $\mathcal{S}^{(m)}$ is computed as in Equation (4) with replacing $e_{g}^{(m)}$ to $\overline{e}_{g}^{(m)}$.
@@
Output: the final threshold $\widehat{t}^{(m)}=t^{(m)}$ for $m\in[M]$ and the derandomized selected set $\mathcal{S}_{\text{derand}}$.
@@
$e_{gr}^{(m)}=G^{(m)}\cdot\frac{\mathbb{I}\left\{g\in\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})\right\}}{\widehat{V}_{\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})}\bigvee\alpha^{(m)}_{0}}$.
@@
where $\{\omega^{(m)}_{r}\}$ denotes the set of normalized weights. We take $\omega^{(m)}_{r}=1/R$ to illustrate the proposed method. Subsequently, by applying the generalized e-filter to the stabilized generalized e-values at the target levels $\alpha^{(1)},\dots,\alpha^{(M)}$, we derive a set of selected features $\mathcal{S}_{\text{derand}}$. Note that for setting (I), $\mathcal{G}_{r}^{(m)}$ is identical to $\mathcal{G}^{(m)}$. And for setting (II), $\mathcal{G}_{r}^{(m)}$ changes when $r$ varies, where weights $\{\omega^{(m)}_{r}\}$ reflect the importance of different procedures.
@@
Given any original vector of FDR level $(\alpha^{(1)}_{0},\dots,\alpha^{(M)}_{0})\in(0,1)^{M}$, any target vector of FDR level $(\alpha^{(1)},\dots,\alpha^{(M)})\in(0,1)^{M}$, and any number of replications $R\geq 1$, for each $m\in[M]$, if $\mathcal{G}^{(m)}\in\mathcal{K}_{\text{finite}}$, then the selected set $\mathcal{S}_{\text{derand}}$ calculated by SFEFP satisfies $\text{FDR}^{(m)}\leq\alpha^{(m)}$ in finite sample settings. Furthermore, if $\mathcal{G}^{(m)}\in\mathcal{K}_{\text{asy}}$ and the group size $|\mathcal{A}_{g}^{(m)}|$ is uniformly bounded, then SFEFP satisfies $\text{FDR}^{(m)}\leq\alpha^{(m)}$ as $(n,N)\to\infty$ at a proper rate.
We further study the stability of SFEFP. We only need to consider setting (I) since the selection set is determined under setting (II). For each layer $m\in[M]$, suppose that the feature or group detection procedure $\mathcal{G}^{(m)}$ is randomized. Conditional on $(\boldsymbol{X},\boldsymbol{y})$, the $e_{gr}^{(m)}$’s are i.i.d. and bounded across each replication $r\in[R]$. By the strong law of large numbers, we have
$\overline{e}_{g}^{(m)}=\frac{1}{R}\sum\limits_{r=1}^{R}e_{gr}^{(m)}\to\bar{\bar{e}}_{g}^{(m)}:=\mathbb{E}[e_{g1}^{(m)}|\boldsymbol{X},\boldsymbol{y}]\quad\text{almost surely as }R\to\infty.$
Let $\mathcal{S}_{\infty}$ denote the set of selected features obtained by SFEFP as $R\to\infty$. The set $\mathcal{S}_{\infty}$ becomes a deterministic function of the data $(\boldsymbol{X},\boldsymbol{y})$ because $\bar{\bar{e}}_{g}^{(m)}$ is fixed. The following theorem demonstrates the stability guaranty of SFEFP.
@@
$\mathbb{P}\left(\mathcal{S}_{\text{derand}}=\mathcal{S}_{\infty}|\boldsymbol{X},\boldsymbol{y}\right)\geq\min\limits_{m\in[M]}\left[1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R}{(G^{(m)})^{2}}\right)\right]^{M}.$
@@ 4 Examples for SFEFP
$\left\{g\in\mathcal{H}_{0}^{\text{grp}}:T_{g}^{(m)}\geq t\right\}\overset{d}{=}\#\left\{g\in\mathcal{H}_{0}^{\text{grp}}:T_{g}^{(m)}\leq-t\right\},\quad{\forall}t>0.$
@@
Consider any $(\alpha^{(1)}_{0},\dots,\alpha^{(M)}_{0})\in(0,1)^{M}$, $(\alpha^{(1)},\dots,\alpha^{(M)})\in(0,1)^{M}$, and any number of replications $R\geq 1$. For each layer $m\in[M]$, suppose that there exists a constant $\tau_{\alpha^{(m)}_{0}}>0$ such that $\mathbb{P}(\text{FDP}^{(m)}(\tau_{\alpha^{(m)}_{0}})\leq\alpha^{(m)}_{0})\to 1$ as $G^{(m)}\to\infty$, the variance $\text{Var}(T_{g}^{(m)})$ is uniformly upper bounded and lower bounded away from zero, $\sum_{g\in\mathcal{H}_{0}^{(m)}}e_{gr}^{(m)}/G^{(m)}$ is uniformly integrable, and $\sum_{g\in\mathcal{H}_{0}^{(m)}}\mathbb{E}[e_{gr}^{(m)}|t^{r}_{\alpha_{0}^{(m)}}>\tau_{\alpha_{0}^{(m)}}]/G^{(m)}$ has a uniform upper bound. If the group size $|\mathcal{A}_{g}^{(m)}|$ is uniformly bounded, then under Assumption 1 and 2, $\{e_{gr}^{(m)}\}_{g\in[G^{(m)}]}$ given in (9) is a set of asymptotic relaxed e-values, and the eDS-filter guarantees that $\limsup_{N\to\infty}\mathrm{FDR}^{(m)}\leq\alpha^{(m)}$.
@@ 5.2 Comparison under low dimensions
Consider $n=1600$, $N=800$, $G=80$, $|\mathcal{H}_{1}|=60$, and $K=20$. Therefore, each group contains 10 features, and on average a relevant group consists of 3 relevant features. The theoretical FDR level is set to $\alpha^{(m)}=0.2$ for $m=1,2$. We first fix the signal strength at $\delta=3$ and vary the correlation $\rho$ in the set $\{0,0.2,0.4,0.6,0.8\}$. Figure 1 illustrates the realized FDR and power at both the individual and group layers as a function of the correlation $\rho$. Next, we fix $\rho=0.7$ and vary $\delta$ in the set $\{3,4,5,6,7\}$. Figure 2 presents these quantities as a function of the signal strength $\delta$, with a fixed correlation $\rho=0.6$. All results are averaged over 50 independent trials.
@@ A.8 Proof of Theorem 5
For each layer $m\in[M]$ and group $g\in[G^{(m)}]$, conditional on the data $(\boldsymbol{X},\boldsymbol{y})$, the $e_{gr}^{(m)}$’s are are independent and identically distributed for different replications $r\in[R]$. By Hoeffding’s inequality, we have
$\begin{split}\mathbb{P}\left(\max\limits_{g\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)&\geq 1-\sum\limits_{g=1}^{G^{(m)}}\mathbb{P}\left(\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|>\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)\\ &\geq 1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R}{(G^{(m)})^{2}}\right).\end{split}$
@@
$\begin{split}\mathbb{P}\left(\bigcap\limits_{m=1}^{M}\left\{\max_{k\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\right\}\Big|\boldsymbol{X},\boldsymbol{y}\right)&=\prod\limits_{m=1}^{M}\mathbb{P}\left(\max_{k\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)\\ &\geq\prod\limits_{m=1}^{M}\left[1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R}{(G^{(m)})^{2}}\right)\right]\\ &\geq\min\limits_{m\in[M]}\left[1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R}{(G^{(m)})^{2}}\right)\right]^{M}.\end{split}$
@@
Thus, the threshold vector $(\widehat{t}^{(1)},\dots,\widehat{t}^{(M)})=(\widehat{t}^{(1)}_{\infty},\dots,\widehat{t}^{(M)}_{\infty})$ belongs to the set $\mathcal{T}(\alpha^{(1)},\dots,\alpha^{(M)})$ and we have $\mathcal{S}_{\infty}\subseteq\mathcal{S}_{\text{derand}}$. Adopting the same argument as above, we also have $\mathcal{S}_{\infty}\supseteq\mathcal{S}_{\text{derand}}$. Hence, the event $\bigcap\limits_{m=1}^{M}\left\{\max_{g\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\right\}$ implies that $\mathcal{S}_{\text{derand}}=\mathcal{S}_{\infty}$. Therefore, we complete the proof that
$\mathbb{P}\left(\mathcal{S}_{\text{derand}}=\mathcal{S}_{\infty}|\boldsymbol{X},\boldsymbol{y}\right)\geq\min\limits_{m\in[M]}\left[1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R}{(G^{(m)})^{2}}\right)\right]^{M}.$

### v5

@@ Abstract
Multiple resolutions arise across a range of explanatory features due to domain-specific structures, leading to the formation of feature groups. It follows that the simultaneous detection of significant features and groups aimed at a specific response with false discovery rate (FDR) control stands as a crucial issue, such as the spatial genome-wide association studies. Nevertheless, existing detection methods with multilayer FDR control generally rely on valid p-values or knockoff statistics, which can be not flexible, powerful and stable in several settings. To fix this issue effectively, this article develops a novel method of Stabilized Flexible E-Filter Procedure (SFEFP), by constructing unified generalized e-values, leveraging a generalized e-filter, and adopting a stabilization treatment with power enhancement. This method flexibly incorporates diverse base detection procedures at different resolutions to provide consistent, powerful, and stable results, while controlling FDR at multiple resolutions simultaneously. Statistical properties of multilayer filtering procedure encompassing one-bit property, multilayer FDR control, and stability guarantee are established. We also develop several examples for SFEFP such as the eDS-filter. Simulation studies and the analysis of HIV mutation data demonstrate the efficacy of SFEFP.
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 rigorously frames the problem of feature detection with multilayer FDR control and introduces necessary preliminaries. Section 3 introduces the proposed SFEFP and relevant theoretical results. In Section 4, we propose eDS-filter and eDS+gKF-filter as examples of SFEFP. Other applications and possible extensions of SFEFP are also discussed. Section 5 presents the performance of our proposed methods on simulated data. In Section 6, the eDS-filter is applied to analyze HIV mutation data. Section 7 discusses the applicable scenarios and the future work of SFEFP. The proofs of all propositions, theorems, and lemmas in the subsequent sections of the paper are provided in the supplementary material.
@@ 3.1 FEFP: Flexible E-Filter Procedure
until $t^{(1)},\ldots,t^{(M)}$ are all unchanged;
@@ 3.2 SFEFP: Stabilized Flexible E-Filter Procedure
for $m=1,\ldots,M$ do
 for $r=1,\dots,R^{(m)}$ do
      Process detection procedure $\mathcal{G}^{(m)}_{r}\in\mathcal{K}_{\text{finite}}\bigcup\mathcal{K}_{\text{asy}}$ with level $\alpha_{0}^{(m)}$.
     Compute the generalized e-value $e_{gr}^{(m)}$ by Equation (6), $g\in[G^{(m)}]$.
@@
  Compute $\overline{e}_{g}^{(m)}=\sum_{r\in[R^{(m)}]}w^{(m)}_{r}e_{gr}^{(m)},g\in[G^{(m)}]$.
end for
@@
     where $\mathcal{S}^{(m)}$ is computed as in Equation (4) with input $\overline{e}_{g}^{(m)}$.
@@
$\widehat{t}^{(m)}\leftarrow t^{(m)}$, $\mathcal{S}_{\text{stab}}=\mathcal{S}(\widehat{t}^{(1)},\dots,\widehat{t}^{(M)})$.
Output: The selected set $\mathcal{S}_{\text{stab}}$.
@@
$e_{gr}^{(m)}=G^{(m)}\cdot\frac{\mathbb{I}\left\{g\in\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})\right\}}{\widehat{V}_{\mathcal{G}^{(m)}_{r}(\alpha_{0}^{(m)})}\bigvee\alpha^{(m)}_{0}}$. (6)
@@
where $\{\omega^{(m)}_{r}\}$ denotes the set of normalized weights. We take $\omega^{(m)}_{r}=1/R$ to illustrate the proposed method. Subsequently, by applying the generalized e-filter to the stabilized generalized e-values at the target levels $\alpha^{(1)},\dots,\alpha^{(M)}$, we derive a set of selected features $\mathcal{S}_{\text{stab}}$. Note that for setting (I), $\mathcal{G}_{r}^{(m)}$ is identical to $\mathcal{G}^{(m)}$. And for setting (II), $\mathcal{G}_{r}^{(m)}$ changes when $r$ varies, where weights $\{\omega^{(m)}_{r}\}$ reflect the importance of different procedures.
@@
Given any original vector of FDR level $(\alpha^{(1)}_{0},\dots,\alpha^{(M)}_{0})\in(0,1)^{M}$, any target vector of FDR level $(\alpha^{(1)},\dots,\alpha^{(M)})\in(0,1)^{M}$, and any number of replications $R\geq 1$, for each $m\in[M]$, if $\mathcal{G}^{(m)}\in\mathcal{K}_{\text{finite}}$, then the selected set $\mathcal{S}_{\text{stab}}$ calculated by SFEFP satisfies $\text{FDR}^{(m)}\leq\alpha^{(m)}$ in finite sample settings. Furthermore, if $\mathcal{G}^{(m)}\in\mathcal{K}_{\text{asy}}$ and the group size $|\mathcal{A}_{g}^{(m)}|$ is uniformly bounded, then SFEFP satisfies $\text{FDR}^{(m)}\leq\alpha^{(m)}$ as $(n,N)\to\infty$ at a proper rate.
We further study the stability of SFEFP. We only need to consider setting (I) since the selection set is determined under setting (II). For each layer $m\in[M]$, suppose that the feature or group detection procedure $\mathcal{G}^{(m)}$ is randomized. Conditional on $(\boldsymbol{X},\boldsymbol{y})$, the $e_{gr}^{(m)}$’s are i.i.d. and bounded across each replication $r\in[R^{(m)}]$. By the strong law of large numbers, we have
$\overline{e}_{g}^{(m)}=\frac{1}{R^{(m)}}\sum\limits_{r=1}^{R^{(m)}}e_{gr}^{(m)}\to\bar{\bar{e}}_{g}^{(m)}:=\mathbb{E}[e_{g1}^{(m)}|\boldsymbol{X},\boldsymbol{y}]\quad\text{almost surely as }R^{(m)}\to\infty.$
Let $\mathcal{S}_{\infty}$ denote the set of selected features obtained by SFEFP as $R^{(m)}\to\infty$, $m\in[M]$. The set $\mathcal{S}_{\infty}$ becomes a deterministic function of the data $(\boldsymbol{X},\boldsymbol{y})$ because $\bar{\bar{e}}_{g}^{(m)}$ is fixed. The following theorem demonstrates the stability guaranty of SFEFP.
@@
$\mathbb{P}\left(\mathcal{S}_{\text{stab}}=\mathcal{S}_{\infty}|\boldsymbol{X},\boldsymbol{y}\right)\geq 1-2\min\left\{\frac{1}{2},\sum_{m=1}^{M}G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R^{(m)}}{(G^{(m)})^{2}}\right)\right\}.$
@@ 4 Examples for SFEFP
$\left\{g\in\mathcal{H}_{0}^{(m)}:T_{g}^{(m)}\geq t\right\}\overset{d}{=}\#\left\{g\in\mathcal{H}_{0}^{(m)}:T_{g}^{(m)}\leq-t\right\},\quad{\forall}t>0.$
@@
Consider any $(\alpha^{(1)}_{0},\dots,\alpha^{(M)}_{0})\in(0,1)^{M}$, $(\alpha^{(1)},\dots,\alpha^{(M)})\in(0,1)^{M}$, and any number of replications $R\geq 1$. For each layer $m\in[M]$, suppose that there exists a constant $\tau_{\alpha^{(m)}_{0}}>0$ such that $\mathbb{P}(\text{FDP}^{(m)}(\tau_{\alpha^{(m)}_{0}})\leq\alpha^{(m)}_{0})\to 1$ as $G^{(m)}\to\infty$, the variance $\text{Var}(T_{g}^{(m)})$ is uniformly upper bounded and lower bounded away from zero, and $\sum_{g\in\mathcal{H}_{0}^{(m)}}e_{gr}^{(m)}/G^{(m)}$ is uniformly integrable. If the group size $|\mathcal{A}_{g}^{(m)}|$ is uniformly bounded, then under Assumption 1 and 2, $\{e_{gr}^{(m)}\}_{g\in[G^{(m)}]}$ given in (10) is a set of asymptotic relaxed e-values, and the eDS-filter guarantees that $\limsup_{N\to\infty}\mathrm{FDR}^{(m)}\leq\alpha^{(m)}$.
@@ 5.2 Comparison under low dimensions
Consider $n=1600$, $N=800$, $G=80$, $|\mathcal{H}_{1}|=60$, and $K=20$. Therefore, each group contains 10 features, and on average a relevant group consists of 3 relevant features. The theoretical FDR level is set to $\alpha^{(m)}=0.2$ for $m=1,2$. We first fix the signal strength at $\delta=3$ and vary the correlation $\rho$ in the set $\{0,0.2,0.4,0.6,0.8\}$. Figure 1 illustrates the realized FDR and power at both the individual and group layers as a function of the correlation $\rho$. Next, we fix $\rho=0.6$ and vary $\delta$ in the set $\{3,4,5,6,7\}$. Figure 2 presents these quantities as a function of the signal strength $\delta$, with a fixed correlation $\rho=0.6$. All results are averaged over 50 independent trials.
@@ A.8 Proof of Theorem 5
For each layer $m\in[M]$ and group $g\in[G^{(m)}]$, conditional on the data $(\boldsymbol{X},\boldsymbol{y})$, the $e_{gr}^{(m)}$’s are are independent and identically distributed for different replications $r\in[R^{(m)}]$. By Hoeffding’s inequality, we have
$\begin{split}\mathbb{P}\left(\max\limits_{g\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)&\geq 1-\sum\limits_{g=1}^{G^{(m)}}\mathbb{P}\left(\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|>\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)\\ &\geq 1-2G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R^{(m)}}{(G^{(m)})^{2}}\right).\end{split}$
@@
$\begin{split}\mathbb{P}\left(\bigcap\limits_{m=1}^{M}\left\{\max_{k\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\right\}\Big|\boldsymbol{X},\boldsymbol{y}\right)&=1-\mathbb{P}\left(\bigcup_{m=1}^{M}\left\{\max_{k\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|\geq\Delta^{(m)}\right\}\big|\boldsymbol{X},\boldsymbol{y}\right)\\ &\geq 1-\sum_{m=1}^{M}\mathbb{P}\left(\max_{k\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|\geq\Delta^{(m)}\big|\boldsymbol{X},\boldsymbol{y}\right)\\ &\geq 1-2\sum_{m=1}^{M}G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R^{(m)}}{(G^{(m)})^{2}}\right).\end{split}$
@@
Thus, the threshold vector $(\widehat{t}^{(1)},\dots,\widehat{t}^{(M)})=(\widehat{t}^{(1)}_{\infty},\dots,\widehat{t}^{(M)}_{\infty})$ belongs to the set $\mathcal{T}(\alpha^{(1)},\dots,\alpha^{(M)})$ and we have $\mathcal{S}_{\infty}\subseteq\mathcal{S}_{\text{stab}}$. Adopting the same argument as above, we also have $\mathcal{S}_{\infty}\supseteq\mathcal{S}_{\text{stab}}$. Hence, the event $\bigcap\limits_{m=1}^{M}\left\{\max_{g\in[G^{(m)}]}\left|\overline{e}_{g}^{(m)}-\bar{\bar{e}}_{g}^{(m)}\right|<\Delta^{(m)}\right\}$ implies that $\mathcal{S}_{\text{stab}}=\mathcal{S}_{\infty}$. Therefore, we complete the proof that
$\mathbb{P}\left(\mathcal{S}_{\text{stab}}=\mathcal{S}_{\infty}|\boldsymbol{X},\boldsymbol{y}\right)\geq 1-2\sum_{m=1}^{M}G^{(m)}\exp\left(\frac{-2(\Delta^{(m)})^{2}R^{(m)}}{(G^{(m)})^{2}}\right).$